Find the general/particular solution of the following Differential Equations


(Non-Exact D.E)


(y - xy³)dx + xdy=0


1
Expert's answer
2021-09-12T18:13:22-0400

Rewrite in the form of a first order Bernoulli ODE


"y'+\dfrac{1}{x}y=y^3"

Substitute "v=y^{1-3}"


"v'=-\dfrac{2}{y^3}y'"

"-\dfrac{1}{2}v'+\dfrac{1}{x}v=1"

Integrating factor "\mu(x)=\dfrac{1}{x^2}"


"\dfrac{1}{x^2}v'-\dfrac{2}{x^3}v=-\dfrac{2}{x^2}"


"(\dfrac{v}{x^2})'=-\dfrac{2}{x^2}"

"\int d(\dfrac{v}{x^2})=-\int\dfrac{2}{x^2}dx"

"\dfrac{v}{x^2}=\dfrac{2}{x}+C"

"v=2x+Cx^2"

Substitute back "v=y^{-2}"


"y^{-2}=2x+Cx^2"

"y^2=\dfrac{1}{2x+Cx^2}"